OpenStudy (anonymous):

a circular coil of radius 5cm and resistance .2ohms is placed in a uniform magnetic field perpendicular to the plane of the coil. the magnitude of the field changes with time according to B=.50e^(-20t)T. what is the magnitude of the current induced in the coil at t=2s?

OpenStudy (anonymous):

This could help: http://hyperphysics.phy-astr.gsu.edu/hbase/magnetic/indcur.html I think that EMF must be calculated first. prior to this flux across coil area must be found. Then, current is easily obtained via Ohm's Law.

OpenStudy (anonymous):

When I wrote flux, I mean derivative of flux relative to time \[d \phi/dt\]

OpenStudy (anonymous):

First of all, find flux using following formula: phi(flux)=B*A.....where B is magnetic field and A is area . using lenz's law: Einduced=-d(phi)/dt find the value of Einduced at t=2. And using ohm's law: I=Einduced/R....Where R is Resistance ...

OpenStudy (anonymous):

i'm not getting the right answer =/... \[emf=d \phi/dt=A*dB/dt\] \[dB/dt=-20*.5e ^{-20t}\]...right? and @ t=2s and A being pi(.05)^2... emf=-3.3*10^-19... I=emf/R=-1.67*10^-18. not even close to any answers

OpenStudy (ujjwal):

i get an answer which is equal to 1.07*10^-18.. is angular velocity= 20 rotations per sec ??

OpenStudy (anonymous):

i didn't use angular velocity

OpenStudy (ujjwal):

I guess you sholuld use angular velocity too